# SubIQ: Inverse Soft-Q Learning for Offline Imitation with Suboptimal Demonstrations

### v1

@@ B.1.1 Mujoco
B.1.2 Panda-gym
The Panda-gym environments (Gallouédec et al., 2021), designed for the Franka Emika Panda robot in reinforcement learning, include PandaReach, PandaPush, PandaPickandPlace, and PandaSlide. Here’s a short description of each:
@@ B.1.2 Panda-gym
B.2 Algorithm Overview
Figure 5: Overview of Sub-IQ.
The Figure 5 provide the overview of our algorithm.
B.3 Qualities of the Generated Expert and Non-Expert Datasets
We create sub-optimal datasets by adding noise to actions of the expert policy and let them interact with the environments to collect the sub-optimal trajectories. Each task has its own difficulty and sensitivity to the noise of the actions. The averaged returns of the generated datasets are reported in Figure 6.
@@ B.3 Qualities of the Generated Expert and Non-Expert Datasets
Ant (5187)
Walker (4477)
Hopper (3748)
Humanoid (8038)
Reach (-1.83)
Push (-3.61)
PnP (-6.61)
Slide (-8.99)
@@
B.4 More details on the Impact of Sub-optimal Datasets
In the Figure 7 we show the evaluation curves with varied sizes for sub-optimal datasets. It can be seem that the overall performance tends to improve with more sup-optimal datasets. In particular, for the Ant task, our sub-IQ even fails to maintain stability when the sizes of sup-optimal datasets are low (10k-5k-1k). Moreover, while BC seems to show improvement with more sub-optimal data, the performance of DWBC and DemoDICE remains unchanged. This may be because these approaches rely on the assumption that expert demonstrations can be extracted from the sub-optimal data, which is not the case in our context.
10k-5k-1k
25k-10k-1k
25k-25k-1k
10k-5k-1k
25k-10k-1k
25k-25k-1k
5k-1k-100
10k-5k-100
10k-10k-100
5k-1k-100
10k-5k-100
10k-10k-100
@@ B.4 More details on the Impact of Sub-optimal Datasets
B.5 Varied Expertise Levels
Figure 8: Average returns of the 5 HalfCheetah datasets (one expert and 4 non-expert).
@@ B.5 Varied Expertise Levels
2 sub-optimal datasets
3 sub-optimal datasets
4 sub-optimal datasets
@@
We assess our algorithm with different numbers of expertise levels to investigate how adding or removing sub-optimal expertise levels influences performance. Specifically, we keep the same expert dataset comprising 1 expert trajectory and conduct tests with 1 to 4 sub-optimal datasets from the Cheetah task. Details of the average returns of the five datasets are reported in Fig. 8. In this context, Sub-IQ outperforms other algorithms in utilizing non-expert demonstrations. Furthermore, BC successfully learns and achieves performance comparable to the performance of sub-IQ with 2 dataset, while DemoDICE and DWBC struggle to learn. The detailed results are plotted in Figure 9. In Section (B.7) below, We particularly delve into the situation of having two datasets (one expert and one sub-optimal), which is a typical setting in prior work.
B.6 Reward Reference Distribution
@@
Ant
Walker
Hopper
Humanoid
Reach
Push
PnP
Slide

### v2

@@ B.3.1 Mujoco
B.3.2 Panda-gym
The Panda-gym environments [12], designed for the Franka Emika Panda robot in reinforcement learning, include PandaReach, PandaPush, PandaPickandPlace, and PandaSlide. Here’s a short description of each:
@@ B.3.2 Panda-gym
B.4 Qualities of the Generated Expert and Non-Expert Datasets
In this paper, we provide a new setting utilizing ranked sub-optimal datasets to perform imitation learning in the offline setting (illustration in Figure 5).
Figure 5: Training datasets illustration.
We create sub-optimal datasets by adding noise to actions of the expert policy and let them interact with the environments to collect the sub-optimal trajectories. Each task has its own difficulty and sensitivity to the noise of the actions. The averaged returns of the generated datasets, computed as percentages w.r.t. the maximum returns, are reported in Figure 6.
@@ B.4 Qualities of the Generated Expert and Non-Expert Datasets
B.5 Hyper Parameters and Experimental Implementations
  • In our experiments, for every algorithm, we run five different seeds corresponding to five different datasets sampled from our databases in Appendix 5.
  • We use double Q critic network for our implementation to increase the stability in the offline training scheme.
  • For IQ-learn [13],SQIL [29], we conduct experiment from its official implement with double Q critic network.
  • For DemoDICE [21] we conduct experiment from its official implementation.
  • For DWBC [37] we conduct experiment from its official implementation.
  • Inspired by double Q-learning, to avoid overfitting in the offline setting, some experiments apply a training trick using KL-divergence between the target actor and the training actor to prevent rapid policy changes.
  • For SAC-based algorithms, we use a fixed exploration parameter which is commonly used in previous work.
  • We conducted all experiments on a total of 8 NVIDIA RTX A5000 GPUs and 64 core CPUs. We use 1 GPUs and 8 core CPUs per task with approximately one day per 5 seeds. The detailed hyper-parameters are reported in Table 2.
Hyper Parameter BC-based SPRINQL
Actor Network [256,256] [256,256]
Critic Network [256,256] [256,256]
Training step 1,000,000 1,000,000
Gamma 0.99 0.99
lr actor 0.0001 0.00003
lr critic 0.0003 0.0003
lr reward reference 0.0003 0.0003
batch size 256 256
soft update critic factor - 0.005
soft update actor factor - 0.00003
exploration temperature - 0.01
reward regularize temperature ($\alpha$) - 1.0
CQL temperature ($\beta$) - 1.0
Table 2: Hyper parameters.
Appendix C Supplementary Experiments
In this section, we present additional experiments that complement those reported in the main paper, along with ablation studies to address the experimental questions stated therein.
C.1 Full Experiment Results for Mujoco and Panda-gym
We report the full results for the 5 different Mujoco environments and 4 different Panda-gym environments, with 3 datasets (i.e., $N=3$), supplementing the results reported in Table 1 in the main paper. The detailed results for Mujoco in Table 3 and Panda-gym in Table 4
Cheetah Ant Walker Hopper Humanoid Avg
BC-E -3.2±$\pm 0.9$ 6.4$\pm$19.1 0.3$\pm$0.4 8.3$\pm$4.1 1.3$\pm$0.2 2.6
BC-O 14.2$\pm$2.9 35.2$\pm$20.1 66.1$\pm$10.8 16.7$\pm$4.0 10.6$\pm$6.3 29.0
BC-both 13.2$\pm$3.6 47.0$\pm$5.9 63.9$\pm$7.3 22.6$\pm$12.8 9.0$\pm$3.5 31.1
W-BC 12.9$\pm$2.8 47.3$\pm$6.4 58.6$\pm$9.1 20.9$\pm$6.6 19.6$\pm$19.0 31.9
TRAIL -4.1$\pm$0.3 -4.7$\pm$1.9 0.2$\pm$0.3 1.1$\pm$0.5 2.6$\pm$0.6 -1.0
IQ-E -3.4$\pm$0.6 -3.4$\pm$1.3 0.1$\pm$0.7 0.3$\pm$0.2 2.4$\pm$0.6 -0.8
IQ-both -6.1$\pm$1.4 -58.2$\pm$0.0 -0.2$\pm$0.1 0.0$\pm$0.0 0.8$\pm$0.0 -12.7
SQIL-E -5.0$\pm$0.7 -33.8$\pm$7.4 0.2$\pm$0.2 0.2$\pm$0.0 0.9$\pm$0.1 -7.5
SQIL-both -5.6$\pm$0.5 -58.0$\pm$0.4 -0.2$\pm$0.0 0.0$\pm$0.0 0.8$\pm$0.0 -12.6
DemoDICE 0.4$\pm$2.0 31.7$\pm$8.9 7.2$\pm$3.1 18.7$\pm$7.5 2.6$\pm$0.8 12.1
DWBC -0.2$\pm$2.5 10.4$\pm$5.0 25.1$\pm$16.3 66.0$\pm$20.9 3.7$\pm$0.3 21.0
SPRINQL (ours) 73.6$\pm$4.3 77.0$\pm$5.6 87.9$\pm$14.2 70.0$\pm$23.8 82.9$\pm$11.2 78.3
Table 3: Comparison results for Mujoco tasks.
Reach Push PnP Slide Avg
BC-E 13.9$\pm$5.9 8.2$\pm$3.8 3.7$\pm$2.7 0.0$\pm$0.0 6.5
BC-O 16.2$\pm$5.5 8.8$\pm$4.5 3.9$\pm$2.7 0.1$\pm$0.3 7.3
BC-both 16.3$\pm$5.0 9.0$\pm$4.3 4.4$\pm$3.0 0.1$\pm$0.4 7.3
W-BC 15.7$\pm$5.1 8.8$\pm$4.3 3.7$\pm$2.8 0.0$\pm$0.0 7.0
TRAIL 18.3$\pm$5.1 11.7$\pm$4.0 7.8$\pm$3.7 1.7$\pm$1.8 9.9
IQ-E 97.7$\pm$2.4 26.3$\pm$10.9 18.1$\pm$12.5 0.1$\pm$0.4 35.6
IQ-both 5.7$\pm$3.4 8.3$\pm$3.9 3.8$\pm$3.3 0.0$\pm$0.2 4.5
SQIL-E 22.1$\pm$15.1 9.6$\pm$3.3 3.2$\pm$2.9 0.1$\pm$0.3 8.8
SQIL-both 8.0$\pm$4.2 8.2$\pm$3.8 3.3$\pm$2.3 0.1$\pm$0.3 4.9
DemoDICE 14.0$\pm$5.3 8.1$\pm$3.7 4.3$\pm$2.4 0.1$\pm$0.5 6.6
DWBC 93.4$\pm$4.3 36.9$\pm$7.4 25.0$\pm$6.3 11.6$\pm$4.4 41.7
SPRINQL (ours) 99.8$\pm$0.9 72.0$\pm$5.3 63.2$\pm$6.4 37.7$\pm$6.6 68.2
Table 4: Comparison results for Panda-gym tasks.
C.2 Comparison Results for $N=2$, i.e., One Expert and One Sub-optimal Datasets - Q1
We provide additional comparison results for $N=2$, complementing to the answer of (Q1), i.e., how does SPRINQL perform compared to other baselines? In this experiment, we want to test the ability of our algorithm in the same scenario of DemoDICE, DWBC which include expert dataset and only one supplementary dataset. Number of expert transitions in Mujoco domains is 1000 and 100 for Panda-gym. Meanwhile, for the supplementary dataset, it is 25000 transitions for Mujoco and 5000 for Panda-gym. In general, although we experience a downgrade in performance due to lack of ranking (only one supplementary dataset), leading to misunderstanding the true reward function, our method is still able to leverage the sub-optimal dataset for understanding the expert demonstrations and provide the highest average score. Reported results are shown in Table 5 6 and Figure 7.
Method Cheetah Ant Walker Hopper Humanoid Avg
BC 3.4$\pm$1.3 39.2$\pm$6.6 45.5$\pm$14.4 28.4$\pm$9.0 24.4$\pm$25.2 28.2
W-BC 3.1$\pm$1.0 39.6$\pm$7.3 46.1$\pm$12.3 29.5$\pm$9.0 28.6$\pm$27.8 29.4
DemoDICE -1.6$\pm$0.6 31.6$\pm$7.5 6.5$\pm$1.7 31.9$\pm$12.3 2.7$\pm$0.6 14.2
DWBC -1.2$\pm$1.3 9.5$\pm$3.4 13.1$\pm$4.7 87.0$\pm$22.2 4.2$\pm$0.4 22.5
SPRINQL (ours) 49.2$\pm$18.5 56.9$\pm$8.2 77.8$\pm$17.6 39.1$\pm$21.0 32.4$\pm$6.4 51.1
Table 5: Comparison results for Mujoco tasks in two expert datasets.
Method Reach Push PnP Slide Avg
BC 17.1$\pm$4.8 8.1$\pm$3.8 3.6$\pm$2.4 0.3$\pm$0.6 7.3
W-BC 18.5$\pm$5.2 9.8$\pm$4.5 3.4$\pm$3.1 0.1$\pm$0.4 8.0
DemoDICE 14.8$\pm$5.7 8.8$\pm$4.4 3.8$\pm$2.9 0.2$\pm$0.6 6.9
DWBC 98.2$\pm$2.4 38.4$\pm$7.4 27.7$\pm$8.1 14.5$\pm$4.3 44.7
SPRINQL (ours) 93.9$\pm$3.3 61.3$\pm$7.2 64.2$\pm$6.7 26.6$\pm$5.5 61.5
Table 6: Comparison results for Panda-gym tasks in two expert datasets.
Cheetah
Ant
Walker
Hopper
Humanoid
Reach
Push
PnP
Slide
Figure 7: Learning curves of two dataset experiment.
C.3 Learning Curves of noReg-SPRINQL and noDM-SPRINQL
This section reports additional results for the comparison between SPRINQL and the two variants noReg-SPRINQL and noDM-SPRINQL (supplementing the answer of Q2 in the main paper). The comparison results for Panda-gym environments are reported in Figure 8 and the learning curves are plotted in Figure 9 and Figure 10.
Reach
Push
PnP
Slide
Figure 8: Ablation study show the performance of three variants of SPRINQL across four Panda-gym environments.
Cheetah
Ant
Walker
Hopper
Humanoid
Reach
Push
PnP
Slide
Figure 9: Two term ablation study for two level dataset scenario.
Cheetah
Ant
Walker
Hopper
Humanoid
Reach
Push
PnP
Slide
Figure 10: Two term ablation study for three level dataset scenario.
C.4 Augmented Expert Demonstrations
We provide experiments to address (Q3) – What happens if we augment the expert data while maintaining the sub-optimal datasets? To this end, we use two Mujoco and two Panda-gym tasks, keeping the same sup-optimal datasets and add more expert demonstrations to the training sets. The comparison results are reported in Figure 11. For the Mujoco tasks, which are more difficult, adding more expert trajectories significantly enhances the performance of all the algorithms. However, for the two Panda-gym tasks, the influence of adding more expert data appears to be less significant in Push and completely absent in PnP. This would be because expert trajectories in these tasks are typically short, consisting of only 2-7 transitions. Hence, a larger quantity of additional expert data may be required to enhance performance.
@@ C.4 Augmented Expert Demonstrations
C.5 Augmented Sup-optimal Demonstrations
Cheetah
Ant
Push
PnP
In this experiment, we want to answer the question (Q4) – What happens if we augment (or reduce) the sub-optimal data while maintaining the expert dataset?. We present numerical results to assess the impact of sub-optimal data on the performance of our SPRINQL and other baselines. To this end, we keep the same expert dataset and adjust the non-expert data used in Table 1. The performance is reported in Fig.12. It is evident that reducing the amount of data in sub-optimal datasets can result in a significant degradation in the performance of our SPRINQL and the other baselines. Conversely, adding more sub-optimal data can enhance overall stability and lead to improved performance.
In the Figure 13 we show the learning curves with varied sizes for sub-optimal datasets. It can be seem that the overall performance tends to improve with more sup-optimal datasets. In particular, for the Ant task, our SPRINQL even fails to maintain stability when the sizes of sup-optimal datasets are low (10k-5k-1k). Moreover, while BC seems to show improvement with more sub-optimal data, the performance of DWBC and DemoDICE remains unchanged. This may be because these approaches rely on the assumption that expert demonstrations can be extracted from the sub-optimal data, which is not the case in our context.
Cheetah
@@ C.5 Augmented Sup-optimal Demonstrations
C.6 Impact of the Conservative Term
In this experiment, we aim to answer (Q5) – How does the conservative term help in our approach? The Equation 8 introduce the conservative Q learning (CQL) term into our work. Here we also test three variants of our method from Section 4.3 and show the impact of CQL to the final performance. The experimental results are shown in Figure 14.
Cheetah
Ant
Push
PnP
Figure 14: Performance of 3 variant with and without the CQL term in four different environments accross two domains.
C.7 Performance with Varying Number of Expertise Levels
Figure 15: Average returns of the 5 HalfCheetah datasets (one expert and four sub-optimals).
@@ C.7 Performance with Varying Number of Expertise Levels
We provide an experiment to answer (Q6) - How does increasing $N$ (the number of expertise levels) affect the performance of SPRINQL? We assess our algorithm with different numbers of expertise levels to investigate how adding or removing sub-optimal expertise levels influences performance. Specifically, we keep the same expert dataset comprising 1 expert trajectory and conduct tests with 1 to 4 sub-optimal datasets from the Cheetah task. Details of the average returns of the five datasets are reported in Fig. 15. In this context, SPRINQL outperforms other algorithms in utilizing non-expert demonstrations. Furthermore, BC successfully learns and achieves performance comparable to the performance of SPRINQL with 2 dataset, while DemoDICE and DWBC struggle to learn. The detailed results are plotted in Figure 16. In Section (C.2) below, We particularly delve into the situation of having two datasets (one expert and one sub-optimal), which is a typical setting in prior work.
C.8 Ablation Study for the Preference-based Weight Learning
We provide this experiment to answer (Q7) – Does the preference-based weight learning approach provide good values for the weights $w_{i}$? To this end, we compare the performance of SPRINQL based on the weights determined by the preference-based methods described in the main paper (denoted as auto W), and the following weighting scenarios:
  • Uniform W: We chose the weights $w_{i}$ uniformly over $[0,1]$ as $\mathbf{w}=\{0.55,0.35,0.15\}$ with a ratio of approximately $10:7:4$.
  • Reduced W: Starting from Uniform W, we reduced the weights of the non-expert data and tested the weights $\mathbf{w}=\{0.65,0.2,0.15\}$ with a ratio of approximately $10:3:2$.
  • Increased W: Starting from Uniform W, we increased the weights of the non-expert data and chose $\mathbf{w}=\{0.4,0.32,0.28\}$ with a ratio of approximately $10:8:7$.
These weight vectors $\mathbf{w}$ are normalized to follow $w_{1}>w_{2}>\ldots>w_{N}$ and $\sum_{i\in[N]}w_{i}=1$.
The comparison results are shown in Figure 17.
Cheetah
Ant
Figure 17: Experiment results for SPRINQL with different manual selection of weight $W$.
C.9 Reward Recovering
In this experiment, we want to answer (Q8) - How does SPRINQL perform in recovering the ground-truth reward function? Compared to BC-based algorithms, one notable advantage of Q-learning based algorithms is their ability to recover the reward function. Here, we present experiments demonstrating reward function recovery across five MuJoCo tasks, comparing recovered rewards to the actual reward function. To achieve this, we introduce increasing levels of random noise to the actions of a trained agent and observe its interactions with the environment. We collect the state, action, and next state for each trajectory, then predict the recovered reward and compare it to the true reward from the environment. For the sake of comparison, we include noReg-SPRINQL, which can be considered an an adaption of IQ-learn [13] to our setting, and noDM-SPRINQL, which is in fact an adaption of T-REX to our offline setting.
Comparison results are presented in Figure 18. We observe a linear relationship between the true and predicted rewards for SPRINQL across all testing tasks, whereas the other approaches fail to return correct relationships for some tasks.
Cheetah
noReg-SPRINQL
noDM-SPRINQL
SPRINQL
Ant
noReg-SPRINQL
noDM-SPRINQL
SPRINQL
Walker
noReg-SPRINQL
noDM-SPRINQL
SPRINQL
Hopper
noReg-SPRINQL
noDM-SPRINQL
SPRINQL
Humanoid
noReg-SPRINQL
noDM-SPRINQL
SPRINQL
Figure 18: Recovered return and the true return of five Mujoco environments.
C.10 Reference Reward Distribution